OpenStudy (anonymous):

WHAT is the EQUATION of the following graph? Already Answered--Need someone to check! Medal will be given!

OpenStudy (anonymous):

can i have two medals?

OpenStudy (anonymous):

you got this

OpenStudy (anonymous):

again by the orientation you see \(x^2\) comes first so it is \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] since it crosses the \(x\) axis at \(2\) you know \(a=2,a^2=9\) and the little box tells you \(b=3,b^2=9\)

OpenStudy (anonymous):

sorry i meant \(a=2,a^2=4\) but anyways you have it